Last year FBGS Inc. had sales of $325,000 and a net income of $19,000, and its year-end assets were $250,000. The firm's total-debt-to-total-capital ratio was 15.0%. The firm finances using only debt and common equity and its total assets equal total invested capital. Based on the DuPont equation, what was the ROE?
Answer #1

total-debt-to-total-capital ratio=Total debt/Total capital

Hence Total debt=(0.15*250,000)=$37500

Total equity=Total capital-Total debt

=(250,000-37500)=$212500

Profit margin=Net income/Sales

=(19000/325000)=0.058461538

Total asset turnover=Sales/Total assets

=(325000/250,000)=1.3

Equity multiplier=Total asset/Equity

=(250,000/212500)=1.176470588

ROE=Profit margin*Total asset turnover*Equity multiplier

=0.058461538*1.3*1.176470588

=8.94%(Approx).